# Particle transport and deposition in wall-sheared thermal turbulence

### v1

@@ 4 Conclusion
In this work, we have performed DNS of wall-sheared thermal turbulence laden with point-particles. We imposed Couette-type shear on the RB convection cell to examine the interactions between vertical buoyancy and horizontal shear forces on particles’ motion. We observed that, with the increase of $Re_{w}$, the large-scale rolls expanded horizontally, evolving into zonal flow in 2-D simulations or streamwise-oriented roll in 3-D simulations. For particles with small and medium $St$, they either circulate within the LSC when buoyancy dominates or drift near the walls when shear dominates. For large $St$, the turbulent flow structure has a minor influence on particles, and they settle almost similarly to those in quiescent fluid environments.
@@
We analyzed particle distribution using Voronoï diagrams and statistical distributions of Voronoï cell areas. For particles with small $St$, the Voronoï cells are randomly distributed and exhibit a low degree of aggregation, which is independent of the flow state being either LSC or zonal flow. For medium $St$, pronounced spatial inhomogeneity and local preferential concentration are observed. For large $St$, clustering also occurs, but wall shears has negligible influence in contrast to those with medium $St$.
@@
We then derive a mathematical model for particle concentration distribution. For particles with very small $St$, we obtain the averaged local particle concentration $\left\langle{{n}\left(y\right)}\right\rangle_{t_{1/2}}/{n_{0}}=1/\left({2\ln 2}\right)$. For very large $St$, we have ${\left\langle{n\left(y\right)}\right\rangle_{t_{1/2}}}/{n_{0}}=2\left({1-y/H}\right)$ at $0.5H\leq y\leq H$ and ${\left\langle{n\left(y\right)}\right\rangle_{t_{1/2}}}/{n_{0}}=1$ at $y\leq 0.5H$. Compared to the DNS results, for a small but nonzero $St$, the local particle concentration generally agrees with the theoretical predictions when the LSC is present, except there is a reduced number of particles in the upper boundary layer; however, when the flow transitions to zonal flow in 2-D simulations, significant deviation appears. For large but not infinite $St$, the local particle concentration generally agrees with theoretical prediction across all $Re_{w}$, highlighting again that the particles’ motion is nearly unaffected by the turbulent flow state.
@@
We further plot the settling curves to quantify deposition ratio. Good agreement is observed between the DNS results and previous theoretical predictions in the small and large $St$ regimes, namely, for small $St$ with an exponential deposition ratio and for large $St$ with a linear deposition ratio. For medium $St$, to bridge the gap, we develop a new model that describes the settling process via an early linear stage and a later non-linear stage. At the early linear settling stage, particles initially released in the lower domain and in the downwelling areas settle. Specifically, the same proportion of particles in the updraft and downdraft regions settle with a speed of $v_{t}-v_{f}$ and $v_{t}+v_{f}$, respectively, leading to the deposition ratio $N_{s}/N_{0}=t/\left({H/{v_{t}}}\right)$. At the later non-linear settling stage, particles circulate within the convection cell during $t_{c}$ and may escape the circulation with a probability of ${\lambda_{f}}$, leading to the deposition ratio $N_{s}/N_{0}=1-{C_{1}}\exp\left[{-{C_{2}}t/\left({H/{v_{t}}}\right)}\right]$. The unknowns in our model can be determined either by least square fitting of the settling curves, or by our empirical relations (29) and (31). In addition, we found that our model for particle deposition is also accurate in predicting the average residence time across a wide range of $St$ for various $Re_{w}$.
@@
Appendix A Spatial distribution of local Kolmogorov length scale.
In the main text, we presented the Kolmogrov length scale based on the volume- and time-averaged method. Here we further present the time-averaged local Kolmogrov length scale distribution $\langle\eta_{K}\rangle_{t}$. The local Kolmogorov length scale is calculated as $\eta_{K}(\mathbf{x},t)=[\nu^{3}/\varepsilon_{u}(\mathbf{x},t)]^{1/4}$ (Shishkina et al., 2010). Taking $Ra=10^{8}$ as an example, we can see from figure 20 that the minimal local Kolmogorov length scale is around 480 $\mu$m, which is six times larger than the largest particle diameter of $d_{p}=80$ $\mu$m. This indicates that the point-particle model can be safely used.
Figure 20: Spatial distribution of time-averaged local Kolmogorov length scale $\langle\eta_{K}\rangle_{t}$ at (a) $Re_{w}=0$, (b) $Re_{w}=2000$, (c) $Re_{w}=4000$, and (d) $Re_{w}=6000$, with $Ra=10^{8}$ and $Pr=0.71$ (see table 1 in the main text for the corresponding fluid properties).
Appendix B Rayleigh number dependence of parameters $t_{s}$ and $C_{2}$.

### v2

@@ 4 Conclusion
In this work, we have performed DNS of wall-sheared thermal turbulence laden with point particles. We imposed Couette-type shear on the RB convection cell to examine the impact of interactions between vertical buoyancy and horizontal shear forces on particles’ motion. We observed that, with the increase of $Re_{w}$, the large-scale rolls expanded horizontally, evolving into zonal flow in 2-D simulations or streamwise-oriented roll in 3-D simulations. For particles with small and medium $St$, they either circulate within the LSC when buoyancy dominates or drift near the walls when shear dominates. For large $St$, the turbulent flow structure has a minor influence on particles, and they settle almost similarly to those in quiescent fluid environments.
@@
We analysed the particle distribution using Voronoï diagrams and statistical distributions of Voronoï cell areas. For particles with small $St$, the Voronoï cells are randomly distributed and exhibit a low degree of aggregation, which is independent of the flow state being either LSC or zonal flow. For medium $St$, pronounced spatial inhomogeneity and local preferential concentration are observed. For large $St$, clustering also occurs, but wall shear has negligible influence in contrast to cases with medium $St$.
@@
We then derive a mathematical model for the particle concentration distribution. For particles with very small $St$, we obtain the averaged local particle concentration $\left\langle{{n}\left(y\right)}\right\rangle_{t_{1/2}}/{n_{0}}=1/\left({2\ln 2}\right)$. For very large $St$, we have ${\left\langle{n\left(y\right)}\right\rangle_{t_{1/2}}}/{n_{0}}=2\left({1-y/H}\right)$ at $0.5H\leq y\leq H$ and ${\left\langle{n\left(y\right)}\right\rangle_{t_{1/2}}}/{n_{0}}=1$ at $y\leq 0.5H$. Compared with the DNS results, for a small but non-zero $St$, the local particle concentration generally agrees with the theoretical predictions when the LSC is present, except there is a reduced number of particles in the upper boundary layer; however, when the flow transitions to zonal flow in 2-D simulations, significant deviation appears. For large but not infinite $St$, the local particle concentration generally agrees with theoretical prediction across all $Re_{w}$, highlighting again that the particles’ motion is nearly unaffected by the turbulent flow state.
@@
We further plot the settling curves to quantify the deposition ratio. Good agreement is observed between the DNS results and previous theoretical predictions in the small and large $St$ regimes, namely, for small $St$ with an exponential deposition ratio and for large $St$ with a linear deposition ratio. For medium $St$, to bridge the gap, we develop a new model that describes the settling process via an early linear stage and a later nonlinear stage. At the early linear settling stage, particles initially released in the lower domain and in the downwelling areas settle. Specifically, the same proportion of particles in the updraft and downdraft regions settle with a speed of $v_{t}-v_{f}$ and $v_{t}+v_{f}$, respectively, leading to the deposition ratio $N_{s}/N_{0}=t/\left({H/{v_{t}}}\right)$. At the later nonlinear settling stage, particles circulate within the convection cell during $t_{c}$ and may escape the circulation with a probability of ${\lambda_{f}}$, leading to the deposition ratio $N_{s}/N_{0}=1-{C_{1}}\exp\left[{-{C_{2}}t/\left({H/{v_{t}}}\right)}\right]$. The unknowns in our model can be determined either by least square fitting of the settling curves, or by our empirical relations (29) and (31). In addition, we found that our model for particle deposition is also accurate in predicting the average residence time across a wide range of $St$ for various $Re_{w}$.
@@
Appendix A Spatial distribution of local Kolmogorov length scale
In the main text, we presented the Kolmogorov length scale based on the volume- and time-averaged method. Here, we further present the time-averaged local Kolmogorov length scale distribution $\langle\eta_{K}\rangle_{t}$. The local Kolmogorov length scale is calculated as $\eta_{K}(\mathbf{x},t)=[\nu^{3}/\varepsilon_{u}(\mathbf{x},t)]^{1/4}$ (Shishkina et al., 2010). Taking $Ra=10^{8}$ as an example, we can see from figure 20 that the minimal local Kolmogorov length scale is around 480 $\mu$m, which is six times larger than the largest particle diameter of $d_{p}=80$ $\mu$m. This indicates that the point-particle model can be safely used.
Figure 20: Spatial distribution of time-averaged local Kolmogorov length scale $\langle\eta_{K}\rangle_{t}$ at (a) $Re_{w}=0$, (b) $Re_{w}=2000$, (c) $Re_{w}=4000$ and (d) $Re_{w}=6000$, with $Ra=10^{8}$ and $Pr=0.71$ (see table 1 in the main text for the corresponding fluid properties).
Appendix B Rayleigh number dependence of parameters $t_{s}$ and $C_{2}$